THE EXPECTED NUMBER OF REAL ROOTS OF A MULTIHOMOGENEOUS SYSTEM OF POLYNOMIAL EQUATIONS

ANDREW MCLENNAN

Abstract. Theorem 1 is a formula expressing the mean number of real roots of a random multihomogeneous system of polynomial equations as a multiple of the mean absolute value of the determinant of a random matrix. Theorem 2 derives closed form expressions for the mean in special cases that include earlier results of Shub and Smale (for the general homogeneous system) and Rojas (for “unmixed” multihomogeneous systems). Theorem 3 gives upper and lower bounds for the mean number of roots, where the lower bound is the square root of the generic number of complex roots, as determined by Bernstein’s theorem. These bounds are derived by induction from recursive inequalities given in Theorem 4.